OpenStudy (anonymous):

sin^2(5) + sin^2(10) +............+sin^2(90)/(1+cos^2(5) + cos^2(10)......cos^2(90)) = ?

OpenStudy (unklerhaukus):

\[\frac{\sin^2(5) + \sin^2(10) +............+\sin^2(90)}{1+\cos^2(5) + \cos^2(10)......\cos^2(90) }= ?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@UnkleRhaukus help

OpenStudy (unklerhaukus):

i can't see a good first step yet. maybe use a trig formula

OpenStudy (unklerhaukus):

are they ment to be more terms in the numerator than the denominator?

OpenStudy (anonymous):

there is 1 in deno

OpenStudy (unklerhaukus):

i mean more in the denominator than the numerator, actually i guess the is an invisible \(\sin(0)\) in the numerator so they have the same number of terms

OpenStudy (unklerhaukus):

the 1 is a really a \(\cos(0)\)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

sorry but there is sin^2(85) in numerator and cos^2(85) in denominator

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

\[=\frac{\sum\limits_{n=0}^{18}\sin(5n)}{\sum\limits_{n=0}^{18}\cos(5n)}\]

OpenStudy (ujjwal):

Add a sin^2 (0) to numerator. And in denominator, 1=cos^2 (0). Now, Sin^2 theta = cos^2 (90-theta) Proceed yourself further.

OpenStudy (experimentx):

sin(5) = cos(85) they are symmetric about 45 ... use these to get 1's

OpenStudy (anonymous):

so wat will we get

OpenStudy (experimentx):

(1 + 1 + 1 + ... +1 + 1/sqrt(2))/(1 + 1 + 1 + ... +1 + 1/sqrt(2)) = 1 my guess ... haven't worked out ... must be 9 1's

OpenStudy (precal):

you are in good hands, can't help you at the moment

OpenStudy (anonymous):

the asnwer should be 1

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

yeah i guessed that ... i also gave you hint how to solve ... it's too long to type..

OpenStudy (ujjwal):

The answer is one. you have already got the process.. Its amazing to see how similar you are to shameer and openstudy1.

OpenStudy (ujjwal):

Try it yourself.

OpenStudy (anonymous):

s i will try

OpenStudy (anonymous):

i got answer as 1/2..........@experimentX

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

you shouldn't get that sin^2(5) + sin^2(10) +............+sin^2(90 = sin^2(0) + sin^2(90) + sin^2(5) + sin^2(85) +............+sin^2(40)+sin^2(50) + sin^2(45) first arrange like this and get 9 + 1/sqrt(2) do same for denominator

OpenStudy (anonymous):

@experimentX 9 + 1/2 for the numerator and the denominator. You forgot to square 1/sqrt(2).

OpenStudy (experimentx):

Oh ... thanks for correcting prof!!

OpenStudy (experimentx):

lol ... silly me!!